# A novel algorithm for optimizing bundle adjustment in image sequence alignment

### v1

@@ 1 Introduction
Electron tomography (ET) techniques is an important branch of cryo-electron microscopy. It utilizes the 3D reconstruction method in simultaneous localization and mapping (SLAM) to reconstruct the 3D structure of organisms by photographing biological samples at different oblique angles. footnotetext: This work was supported by the Original Exploratory Program Project of National Natural Science Foundation of China (62450004), the Joint Funds of the National Natural Science Foundation of China (U23A20325), and the Natural Science Foundation of Shandong Province (ZR2021ZD14, ZR2024MF045). (Corresponding author: Huanshui Zhang.) H. Zhang is with the College of Electrical Engineering and Automation, Shandong University of Science and Technology, Qingdao 266590, China, and also with the School of Control Science and Engineering, Shandong University, Jinan, Shandong 250061, China (e-mail: hszhang@sdu.edu.cn). H. Wang is with the College of Electrical Engineering and Automation, Shandong University of Science and Technology, Qingdao 266590, China (e-mail: whx1123@126.com;). H. Xu is with the College of Mathematics and Systems Science, Shandong University of Science and Technology, Qingdao 266590, China (e-mail: 201901070720@sdu.edu.cn;)This method enables the acquisition of relatively accurate biological structural information. The most critical step is the alignment of image sequence, which has a direct impact on the final 3D structure restoration accuracy. Image sequence alignment can be divided into two main methods: marker-free alignment [1] and marker-based alignment [2, 3]. In the marker-based alignment , the bundle adjustment is the core, and the nonlinear least square method is usually used to minimize the reprojection error and optimize the target parameters.
@@ 2.1 Levenberg-Marquardt algorithm
Let $\varphi(x)=(\varphi_{1}(x),\varphi_{2}(x),...,\varphi_{m}(x))^{T}$, the objective is to minimize the sum of squared residuals
@@
$\left(J^{T}(x_{k})J(x_{k})+\mu_{k}I\right)d_{k}$ $=-J^{T}(x_{k})\varphi(x_{k}),$ (2)
@@
$\left(J^{T}(x_{k})J(x_{k})+\mu_{k}I\right)d_{k}=-J^{T}(x_{k})\varphi(x_{k})$
@@ 3.1 Projection model
where $(X,Y,Z)^{T}$ is the 3D coordinates of the fiducial markers, $(u,v)^{T}$ is the 2D projection of the fiducial markers in the image, $s$ is the image scale change, $\textbf{t}=(t_{0},t_{1})^{T}$ is the translation during projection and P represents the orthogonal projection matrix. R denotes the rotation matrix of the projection process, the details of R and P as follows
@@
$\textbf{x}_{c}=(s,\alpha,\beta,\gamma,t_{0},t_{1})^{T}$ is called as camera parameters and $\textbf{x}_{p}=(X,Y,Z)^{T}$ is called as point parameters.
@@
Let the observed point be denoted as $\textbf{z}\stackrel{{\scriptstyle\Delta}}{{=}}(u,v)^{T}$ and the point obtained by reprojection can be expressed as follows
@@ 3.2 Optimal control algorithm
Suppose $f(x):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a nonlinear function with a continuous second derivative. Consider the following optimization problem
$\underset{x\in\mathbb{R}^{n}}{min}\quad f(x).$ (11)
Generally it can be solved by the following iteration
@@
where $d_{k}$ is step size.
If we interpret $d_{k}$ as a control input in a control system, the iteration (12) exhibits a structure analogous to that of the following discrete-time linear time-invariant system
@@
By designing a performance index related to the cost function [4], the optimization problem (11) can be reformulated as the following optimal control problem,
@@
Based on the Pontryagin Maximum Principle, the optimal trajectory satisfies a set of forward and backward equations. Solving these equations yields the optimal controller for the problem (14)
$u_{k}=-R^{-1}\sum_{i=k+1}^{N+1}\nabla f(x_{i}).$ (15)
Thus, we can derive the following iterative algorithm
@@
Although the above iteration is implicit, we can obtain an approximate explicit solution through Taylor expansion. First, take the first-order Taylor expansion of $\nabla f(x_{i})$ at $x_{i-1}$
$x_{k+1}=x_{k}-g_{k},k=0,\dots,N,$ (17)
$g_{k}=[R+H_{k}]^{-1}[\nabla f_{k}+\sum_{i=k+1}^{N}(\nabla f_{i}-H_{i}g_{i})],$ (18)
$g_{N}=[R+H_{k}]^{-1}\nabla f_{N},$ (19)
where $\nabla f_{k}=\nabla f(x_{k}),H_{k}=H(x_{k})$, and $H(x_{k})$ is the Hessian matrix of the objective function $f(x)$. By replacing all $x_{i}$ with $x_{k}$ for $i>k$, we obtain the following optimal control algorithm to address BA
$x_{k+1}=x_{k}-g_{k}(x_{k}),$
$g_{0}(x_{k})=[R+H(x_{k})]^{-1}\nabla f(x_{k}),$ (20)
$g_{k}(x_{k})=[R+H(x_{k})]^{-1}[\nabla f(x_{k})+Rg_{k-1}(x_{k})].$
The optimal control algorithm has been demonstrated to exhibit a superlinear convergence property[11], and the algorithm operates through the following stages
@@ 4.1 Real datasets experiment
Data Method R/$\mu_{0}$ Iteration Initial residual Final residual
Centriole OCA $R=0.08I$ 4 1.729723 1.196488
L-M $\mu_{0}=0.1$ 9 1.196488
VEEV OCA $R=17I$ 28 10.412596 1.202426
L-M $\mu_{0}=0.1$ 236 1.202426
Vibrio OCA $R=0.01I$ 5 10.431754 1.123689
L-M $\mu_{0}=0.1$ 7 1.123689
@@ 4.2 Simulated datasets experiment
To obtain a broader range of experimental results, we randomly constructed $n$ 3D points and $m$ projection images at various angles for simulated dataset experiments, where $n=20,40,60$ and $m=21,41,64$. The projection images are sized $1024\times 1024$ pixels, and the 3D points are distributed within an $800\times 800\times 400$ space. To enhance the realism of the simulated dataset, we processed the constructed data as follows
@@
For convenience, we denote the dataset with parameters $m=21,n=20,a=0.2,b=5$ as “21c_5%_20p_0.2%.” To ensure a fair comparison, the weight matrix $R$ in the OCA and the damping factor $\mu$ in the L-M algorithm were iteratively adjusted to promote faster convergence. The results of this experiment are presented in Table 2.
@@
Noise Method $R/\mu_{0}$ Iteration Initial residual Final residual
21c_5OCA $R=0.25I$ 6 11.178429 2.051380
_20p_0.2L-M $\mu_{0}=0.1$ 8 2.051380
21c_5OCA $R=0.25I$ 6 14.683220 12.377970
_20p_2L-M $\mu_{0}=1$ 24 12.377970
41c_5OCA $R=0.25I$ 5 8.411878 1.920106
_20p_0.2L-M $\mu_{0}=0.1$ 6 1.920106
41c_5OCA $R=0.625I$ 5 15.658034 12.106996
_20p_2L-M $\mu_{0}=10$ 17 12.106996
21c_5OCA $R=0.25I$ 6 8.085499 1.992593
_40p_0.2L-M $\mu_{0}=0.1$ 6 1.992593
21c_5OCA $R=I$ 8 14.191696 11.834401
_40p_2L-M $\mu_{0}=1$ 16 11.834401
41c_5OCA $R=0.25I$ 5 7.968274 1.860539
_40p_0.2L-M $\mu_{0}=0.1$ 6 1.860539
41c_5OCA $R=0.5I$ 5 16.647453 13.511510
_40p_2L-M $\mu_{0}=1$ 9 13.511510
When large noise is introduced into the simulated dataset, the initial estimates of the camera parameters and the 2D point coordinates exhibit significant deviations. This suggests that the initial parameters are considerably distant from the extremum, necessitating a greater number of iterations for both the OCA and L-M algorithm. It is worth noting that, in the L-M algorithm, the damping factor $\mu$ is continuously updated, whereas the weight matrix $R$ in the OCA remains constant throughout the iterations. If we allow the matrix $R$ in the OCA to be updated gradually similar to $\mu$ in the L-M algorithm, the OCA could potentially achieve faster convergence. The detailed experimental results are displayed in the Table 3.
@@
Noise Method $R_{0}/\mu_{0}$ Iteration Initial residual Final residual
21c_10OCA $R_{0}=10^{5}I$ 8 75.677900 67.390553
_20p_10L-M $\mu_{0}=0.01$ 108 67.390553
41c_10OCA $R_{0}=2*10^{5}I$ 9 68.691760 58.815270
_20p_10L-M $\mu_{0}=0.1$ 146 58.815270
21c_10OCA $R_{0}=75I$ 8 70.903339 65.472015
_40p_10L-M $\mu_{0}=100$ 168 65.472015
41c_10OCA $R_{0}=10I$ 8 79.020956 73.499860
_40p_10L-M $\mu_{0}=0.1$ 26 73.499860

### v2

@@ 1 Introduction
Electron tomography (ET) techniques is an important branch of cryo-electron microscopy. It utilizes the 3D reconstruction method in simultaneous localization and mapping (SLAM) to reconstruct the 3D structure of organisms by photographing biological samples at different oblique angles. footnotetext: This work was supported by the Original Exploratory Program Project of National Natural Science Foundation of China (62450004), the Joint Funds of the National Natural Science Foundation of China (U23A20325), and the Natural Science Foundation of Shandong Province (ZR2021ZD14, ZR2024MF045). (Corresponding author: Renmin Han, Huanshui Zhang.) H. Zhang is with the College of Electrical Engineering and Automation, Shandong University of Science and Technology, Qingdao 266590, China, and also with the School of Control Science and Engineering, Shandong University, Jinan, Shandong 250061, China (e-mail: hszhang@sdu.edu.cn). R. Han is with the Research Center for Mathematics and interdisciplinary Sciences, Shandong University, Qingdao 266237, China. (e-mail: hanrenmin@sdu.edu.cn). H. Wang is with the College of Electrical Engineering and Automation, Shandong University of Science and Technology, Qingdao 266590, China (e-mail: whx1123@126.com;). H. Xu is with the College of Mathematics and Systems Science, Shandong University of Science and Technology, Qingdao 266590, China (e-mail: 201901070720@sdu.edu.cn;)This method enables the acquisition of relatively accurate biological structural information. The most critical step is the alignment of image sequence, which has a direct impact on the final 3D structure restoration accuracy. Image sequence alignment can be divided into two main methods: marker-free alignment [1] and marker-based alignment [2, 3]. In the marker-based alignment , the bundle adjustment is the core, and the nonlinear least square method is usually used to minimize the reprojection error and optimize the target parameters.
@@ 2.1 Levenberg-Marquardt algorithm
Let $\varphi(x)=(\varphi_{1}(x),\varphi_{2}(x),...,\varphi_{m}(x))^{\top}$, the objective is to minimize the sum of squared residuals
@@
$\left(J^{\top}(x_{k})J(x_{k})+\mu_{k}I\right)d_{k}$ $=-J^{\top}(x_{k})\varphi(x_{k}),$ (2)
@@
$\left(J^{\top}(x_{k})J(x_{k})+\mu_{k}I\right)d_{k}=-J^{\top}(x_{k})\varphi(x_{k})$
@@ 3.1 Projection model
where $(X,Y,Z)^{T}$ is the 3D coordinates of the fiducial markers, $(u,v)^{\top}$ is the 2D projection of the fiducial markers in the image, $s$ is the image scale change, $\textbf{t}=(t_{0},t_{1})^{\top}$ is the translation during projection and P represents the orthogonal projection matrix. R denotes the rotation matrix of the projection process, the details of R and P as follows
@@
$\textbf{x}_{c}=(s,\alpha,\beta,\gamma,t_{0},t_{1})^{\top}$ is called as camera parameters and $\textbf{x}_{p}=(X,Y,Z)^{\top}$ is called as point parameters.
@@
Let the observed point be denoted as $\textbf{z}\stackrel{{\scriptstyle\Delta}}{{=}}(u,v)^{\top}$ and the point obtained by reprojection can be expressed as follows
@@ 3.2 Optimal control algorithm
Suppose $f(x):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a nonlinear function with a continuous second derivative. Consider the following optimization problem:
$\underset{x\in\mathbb{R}^{n}}{\text{min}}\quad f(x).$ (11)
Generally, it can be solved by the following iteration:
@@
where $d_{k}$ is the step size.
If we interpret $d_{k}$ as a control input in a control system, the iteration (12) exhibits a structure analogous to that of the following discrete-time linear time-invariant system:
@@
By designing a performance index related to the cost function, the optimization problem (11) can be reformulated as the following optimal control problem:
@@
Based on the Pontryagin Maximum Principle, the optimal trajectory satisfies a set of forward and backward equations. Solving these equations yields the optimal controller for the problem (14):
$d_{k}=-R^{-1}\sum_{i=k+1}^{N+1}\nabla f(x_{i}).$ (15)
Thus, we can derive the following iterative algorithm:
@@
Although the above iteration is implicit, it can be obtained as an approximate explicit solution through Taylor expansion. First, take the first-order Taylor expansion of $\nabla f(x_{i})$ at $x_{i-1}$:
$x_{k+1}=x_{k}-d_{k}(x_{k}),\quad k=0,...,N,$ (17)
$d_{N}(x_{k})=[R+H(x_{k})]^{-1}f(x_{N}),$ (18)
$d_{l}(x_{k})=[R+H(x_{k})]^{-1}[\nabla f(x_{k})+\sum_{i=k+1}^{N}(\nabla f(x_{i})-H(x_{i})d_{i}(x_{k}))],\quad l=N-1,\dots,k,$ (19)
where $H(x)$ is the Hessian matrix of $f(x)$. By substituting $x_{k}$ for all $x_{i}(i>k)$ on the right-hand side of equation (19) and transforming the backward iteration scheme into a forward formulation, we derive the following optimal control algorithm to solve the BA model:
$x_{k+1}=x_{k}-d_{k}(x_{k}),\quad k=0,...,N,$
$d_{0}(x_{k})=[R+H(x_{k})]^{-1}\nabla f(x_{k}),$ (20)
$d_{l}(x_{k})=[R+H(x_{k})]^{-1}[\nabla f(x_{k})+Rd_{l-1}(x_{k})],\quad l=1,\dots,k$
The optimal control algorithm has been demonstrated to exhibit a superlinear convergence and operates through the following stages. The details can be referred to [4].
@@ 4.1 Real datasets experiment
{tblr}
colspec = c c c c c c c, vlines, hlines, row1 = font=, cell2,4,6,81 = r=2c, cell2,4,6,85 = r=2c, Data & Method $R/\mu_{0}$ Iteration Initial residual Final residual Time
Centriole OCA $R=0.08I$ 4 1.730 1.196 0.068s
L-M $\mu_{0}=0.1$ 9 1.196 0.193s
VEEV OCA $R=17I$ 28 10.413 1.202 0.176s
L-M $\mu_{0}=0.1$ 236 1.202 1.776s
Vibrio OCA $R=0.01I$ 5 10.431 1.124 0.372s
[…]
Adhesion belt OCA $R=5I$ 12 1.065 0.477 0.751s
[…]
@@ 4.2 Simulated datasets experiment
To obtain a broader range of experimental results, we randomly constructed $n$ 3D points and $m$ projection images at various angles for simulated dataset experiments, where $n=20,40,60$ and $m=21,41,64$. The projection images are sized $1024\times 1024$ pixels, and the 3D points are distributed within an $800\times 800\times 400$ space. To enhance the realism of the simulated dataset, we processed the constructed data as follows:
@@
For convenience, we denote the dataset with parameters $m=21,n=20,a=0.2,b=5$ as “m21 n20 a0.2 b5”. To ensure a fair comparison, the weight matrix $R$ in the OCA and the damping factor $\mu$ in the L-M algorithm were iteratively adjusted to promote faster convergence. The results of this experiment are presented in Table 2.
@@
{tblr}
width = colspec = c c c c c c c, vlines, hlines, row1 = font=, cell2,4,6,8,10,12,14,161 = c=1, cell2,4,6,8,10,12,14,165 = c=1, Data & Method $R/\mu_{0}$ Iteration Initial residual Final residual Time
[…]
L-M $\mu_{0}=0.1$ 8 2.051 0.032s
[…]
L-M $\mu_{0}=1$ 24 12.378 0.090s
[…]
L-M $\mu_{0}=0.1$ 6 1.920 0.050s
[…]
L-M $\mu_{0}=10$ 17 12.107 0.144s
[…]
L-M $\mu_{0}=0.1$ 6 1.992 0.042s
[…]
L-M $\mu_{0}=1$ 16 11.834 0.010s
[…]
L-M $\mu_{0}=0.1$ 6 1.861 0.089s
[…]
L-M $\mu_{0}=1$ 9 13.512 0.153s
When the high noise is introduced into the simulated dataset, the initial estimates of the camera parameters and the 2D point coordinates exhibit significant deviations. This suggests that the initial parameters are considerably distant from the extremum, necessitating a greater number of iterations for both the OCA and L-M algorithm. It is worth noting that, in the L-M algorithm, the damping factor $\mu$ is continuously updated, whereas the weight matrix $R$ in the OCA remains constant throughout the iterations. If we allow the matrix $R$ in the OCA to be updated gradually similar to $\mu$ in the L-M algorithm, the OCA could potentially achieve faster convergence. The detailed experimental results are displayed in Table 3.
@@
{tblr}
width = colspec = c c c c c c c, vlines, hlines, row1 = font=, cell2,4,6,81 = c=1, cell2,4,6,85 = c=1, Data & Method $R_{0}/\mu_{0}$ Iteration Initial residual Final residual Time
[…]
L-M $\mu_{0}=0.01$ 108 67.391 0.411s
[…]
L-M $\mu_{0}=0.1$ 146 58.815 0.100s
[…]
L-M $\mu_{0}=100$ 168 65.472 1.133s
[…]
L-M $\mu_{0}=0.1$ 30 57.466 0.452s